OpenStudy (anonymous):

(w+t)/(wt^2 )+(9w+t)/(w^2 t)

OpenStudy (anonymous):

w^2+10tw+t^2

myininaya (myininaya):

\[\frac{(w+t)*w}{wt^2*w}+\frac{(9w+t)*t}{w^2t*t}\]

myininaya (myininaya):

\[\frac{w^2+tw+9wt+t^2}{w^2t^2}=\frac{w^2+10wt+t^2}{w^2t^2}\]

OpenStudy (anonymous):

w^2+10tw+t^2 / w^2t^2 Forgot about that. Sorry.